OpenStudy (anonymous):

Find the points of inflection for the y=x^5-15x^3

OpenStudy (anonymous):

take the derivative twice. set the result = 0 and solve

OpenStudy (anonymous):

y'=5x^4-45x^2 y''=20x^3-90x 0 =10x(2x^2-9) \[x=0, x=\pm \sqrt{\frac{9}{2}}\] so are my intervals going to be

OpenStudy (anonymous):

\[y=x^5-15x^3\] \[y'=5x^4-45x^2\] \[y''=20x^3-90x\]

OpenStudy (anonymous):

oh looks like you got it yes?

OpenStudy (anonymous):

\[10x(2x^2-9)=0\] \[x=0\] \[x=\pm\sqrt{\frac{9}{2}}\]

OpenStudy (anonymous):

forget sign chart

OpenStudy (anonymous):

i cant i have to do it that way...lol

OpenStudy (anonymous):

you have a third degree polynomial with positive leading coefficient yes?

OpenStudy (anonymous):

hell no. think of what the picture looks like. you have 3 zeros \[a,b,c\] and a third degree polynomial. it will be negative on \[(\infty,a)\]

OpenStudy (anonymous):

positive on \[(a,b)\] negative of \[(b,c)\] and positive on \[(c,\infty)\]

OpenStudy (anonymous):

Do i use all those intervals in my sign chart \[(-\infty,-\sqrt{\frac{9}{2}}), (-\sqrt{\frac{9}{2}},0), (0,\sqrt{\frac{9}{2}}), (\sqrt{\frac{9}{2}},\infty)\]

OpenStudy (anonymous):

yes and it will be negative positive negative positive respectively

OpenStudy (anonymous):

i have to use all of 'em?

OpenStudy (anonymous):

making your original 5th degree poly concave down concave up concave down concave up

OpenStudy (anonymous):

yes you have 4 intervals

OpenStudy (anonymous):

lol okay thanks satellite!!!

OpenStudy (anonymous):

yw. btw throw out that "sign chart" it is a time waster. just remember what a third degree poly looks like and use common sense

OpenStudy (anonymous):

for the 2nd interval i got a negative

OpenStudy (anonymous):

my lecturer says we have to use it lol...

OpenStudy (anonymous):

oh nvm no i didnt

OpenStudy (anonymous):

tell him/her not to waste your time. why do they bother to teach what a third degree poly looks like if you cannot use it later?

OpenStudy (anonymous):

lol idk

OpenStudy (anonymous):

what a nice picture http://www2.wolframalpha.com/input/?i=y%3D20x^3-90x

OpenStudy (anonymous):

as you can see.... negative positive negative positive sign chart be damned

OpenStudy (anonymous):

Refer to the attachment

OpenStudy (anonymous):

okay thanks satellite

OpenStudy (anonymous):

oh wow thanks robtobey!

OpenStudy (anonymous):

Your welcome.